1/2

N. N. 42,

The House of the Rising Sun

Versio de la brita muzik-grupo The Animals

There is a house in New Orleans They call the Rising Sun And it’s been the ruin of many a poor boy And God I know I’m one.

My mother was a tailor She sewed my new bluejeans My father was a gamblin’ man Down in New Orleans

Now the only thing a gambler needs Is a suitcase and trunk And the only time he’s satisfied Is when he’s on a drunk

Oh mother tell your children Not to do what I have done Spend your lives in sin and misery In the House of the Rising Sun

Well, I got one foot on the platform The other foot on the train I’m goin’ back to New Orleans To wear that ball and chain

…

2/2

Well, there is a house in New Orleans They call the Rising Sun And it’s been the ruin of many a poor boy And God I know I’m one.

Verkinto de tiu ĉi Angla poemo estas N. N. 42.

Arg-719-1444 (2012-12-10 11:04:10)

Vidu la retejojn http: // de. wikipedia. org/ wiki/ The_House_of_the_Rising_Sun kaj http: // de. wikipedia. org/ wiki/ The_Animals .

http://www.poezio.net/version?poem-id=719&version-id=,1444